Last year Maria invested $10,000, part at 6% annual interest and the rest at 8% annual interest. if she received $760 in interest at the end of the year, how much dis she invest at each rate?
Answer by stanbon(75887)

Last year Maria invested $10,000, part at 6% annual interest and the rest at 8% annual interest. if she received $760 in interest at the end of the year, how much dis she invest at each rate?
-----
Quantity Eq: x + y = 10,000
Interest Eq:0.06x+0.08y = 760
-----------------------------------
Multiply thru 1st Eq. by 6
Multiply thru 2nd Eq. by 100
-------
6x + 6y = 6*10000
6x + 8y = 76000
---------------
Subtract 1st equation from 2nd and solve for "y":
2y = 16000
y = $8000 (amt. invested at 8%)
----
Solve for "y":
x + y = 10,000
x + 8000 = 10,000
x = $2000 (amt. invested at 6%)
=================================
Cheers,
Stan H.
==========